OpenStudy (anonymous):

[3f(x)+f(x)g(x)]=36 find f(2) suppose f and g are continuous functions such that g(2)=6

OpenStudy (campbell_st):

ok... start by substituting g(2) = 6 then you have 3f(2) + 6f(2) = 36 solve for f(2)

OpenStudy (anonymous):

Thanks, so it it will be 9f(2)= 36, f(2)= 4. My question is, how can you easily just plug in? does f(x)= f(2) because they're continuous and lim x->a f(x)= f(a) so g(2)= f(2)?

OpenStudy (campbell_st):

thats whta I got... and it did ask to find f(2)

OpenStudy (anonymous):

kk thank you i get it